The flat rate of taxis is 45 pesos. For every half-kilometer,
2 pesos and 50 centavos are added to the fare. What is the
equation for the total fare if k is the distance traveled in
kilometers and t is the cost?
Thank you in advance!

Answer by Theo(13342)

the flat rate is 45 pesos.
there is an additional charge of 2 pesos and 50 centavos for each half kilometer traveled.

t is the total cost.
k is the number of kilometers traveled.
2 * k is the number of half kilometers traveled.

there are 100 centavos to 1 peso.
therefore 50 centavos is equal to .5 pesos.
this means the additional charge for each half-kilometer is 2.5 pesos.
this means the additional charge for each kilometer is 5 pesos.

the formula becomes t = 45 + 5 * k
t is the total cost.
k is the number of kilometers.

without any rounding or truncating of kilometers, the formula would work like this.

you get in the taxi and go for 20.5 kilometers.
the total charge is 45 + 5 * 20.5 = 147.5 pesos.
that would be 147 pesos and 50 centavos.

using another example, the formula would work like this.

you get in the taxi and go for 20.5436 kilometers.
the total charge is 45 + 5 * 20.5436 = 147.718 pesos.
you would round to the nearest centavo.
the total charge becomes 147.72 pesos.
that would be 147 pesos and 72 centavos.

if you round to the nearest half kilometer, then the total charges is 45 + 5 * 20.5 = 147.5 pesos.
that would be 147 pesos and 50 centavos.

the formula is good without taking rounding into consideration.
the rounding is based on the policy of the taxi company as well as the accuracy of the odometer.
the cost would probably be rounded to the nearest centavo.
the kilometers could be rounded up or rounded down, or just rounded to the nearest half kilometer as in the last example.

an example of rounding to the nearest half kilometer follows.

20.5436 is rounded to 20.5 because it's closer to 20.5 than it is to 21.
20.5436 is rounded up to 21
20.5436 is rounded down to 20.5

20.9803 is rounded to 21 because it's closer to 21 than it is to 20.5.
20.9803 is rounded up to 21
20.9803 is rounded down to 20.5